OpenStudy (anonymous):

If 5(2+4z)=85,what does z=?

OpenStudy (anonymous):

5(2+4z)=85 10+20z=85 20z=75 z=15/4

jimthompson5910 (jim_thompson5910):

5(2+4z)=85 10+20z=85 20z = 85-10 20z = 75 z = 75/20 z = 15/4

OpenStudy (anonymous):

5(2+4z)=85 2+4z=85/5 2+4z=17 4z=15 z=15/4

OpenStudy (anonymous):

15/4